# Error thresholds of toric codes with transversal logical gates

### v2

@@ III.1 Persistent bit-flip error model
Therefore, we consider an error model that captures the spreading of error, where bit-flip errors occur both before and after the tCNOT gate. To directly compare with the case of quantum memory, we evenly split the error rate of channel $\mathcal{N}_{p}$ into two bit-flip channels $\mathcal{N}_{\tilde{p}}$ before and after the tCNOT gate:
@@
We will subsequently refer to such an error model as the persistent bit-flip error [^3] When referring to the bit-flip error without the word “persistent”, we assume the bit-flip channel in Eq. (6) occurs only after the tCNOT gate, which is the case in the noise model we study in Sec. IV. .
@@
III.2 Detectors of toric codes across a tCNOT gate
The main challenge in applying stat-mech mapping to logical circuits lies in the necessity of tracking dynamical error propagation through gates. For this, we import the quantum circuit simulation notion of detector [41], and integrate it into the stat-mech mapping procedure. In our case, a detector associated with a stabilizer measurement outcome is a product of it with earlier stabilizer measurement outcomes that back-tracks the flow of that measured Pauli operator in the circuit. Without errors, the value of each detector is deterministically 1. When an error $E$ occurs in the circuit, a detector can be triggered, which flips its value to -1. The collection of detector values, $\{d\}$, serves as the syndrome. Therefore, once the detectors and its relation with every error $E$ in the error model is specified, so will the cycles $C$.
In the context of a logical circuit with a tCNOT gate, the spacetime detectors are fixed in the following way [24, 26, 27, 28, 29]. Since we only consider bit-flip errors, we back-track the spread of the measured $Z$ stabilizers. In this way, the spacetime detectors between the two adjacent syndrome extraction rounds are the ones shown in Fig. 2(a): the spacetime detector $d^{c}$ of the control code block is the product of the measurement outcomes $M^{c}_{1/2}$ of the $Z$ stabilizers before/after the tCNOT gate:
@@ III.2 Detectors of toric codes across a tCNOT gate
where $M^{t}_{1,2}$ are outcomes of the $Z$ stabilizer measurements before and after the tCNOT gate. The superscript denotes the control code ($c$) and the target code ($t$) respectively (alternatively, one can also choose $d^{t}=M^{c}_{2}M^{t}_{1}M^{t}_{2}$.). A visualization of the detectors is show in Fig. 2(a). The spacetime detectors directly incorporate both error propagation mechanisms caused by the tCNOT gate. Namely, a bit-flip error that occurs in the error channel $\mathcal{N}_{\tilde{p}}$ before the tCNOT gate in the control block now triggers both $d^{c}$ and $d^{t}$.
Figure 2: (a) The implementation of the tCNOT gate between two rounds of syndrome extractions. The black dots at the centers of shaded plaquettes represent the measurement of the $Z$ stabilizers defined on these plaquettes. The choices of spacetime detectors $d^{c}$ and $d^{t}$ for the control and target code blocks between the two rounds of SE are marked in blue and green, repectively. Due to the tCNOT gate, the spacetime detector of the target block, $d^{t}$, extends to the control block, whose value is obtained from multiplying three $Z$ stabilizer measurement outcomes $d^{t}=M^{c}_{1}M^{t}_{1}M^{t}_{2}$. To avoid visual overlap, we sketch the spacetime detectors $d^{c}$ and $d^{t}$ in two adjacent plaquettes that host $Z$ stabilizers. (b) The logical circuit with tCNOT gate, persistent bit-flip noise channels $\mathcal{N}_{\tilde{p}}$ and two rounds of perfect syndrome extractions (SE), during which the weight-4 $Z$ stabilizers are measured.
III.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
With the detectors and error model fixed, we now construct the stat-mech model of the two toric code blocks across the tCNOT gate. We denote the bit-flip errors that occur during the four error channels $\mathcal{N}_{\tilde{p}}$ in Fig. 2(a) as $E^{c,t}_{1,2}$, where $c/t$ indicates the control/target code blocks and $1/2$ denotes the error before/after the tCNOT gate. Given the form of the spacetime detectors $d^{c/t}$, the errors that trigger spacetime detectors in the control and target blocks are
@@ III.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
We focus on a pair of physical qubits in the control and target block. Since the error channels $\mathcal{N}_{\tilde{p}}$ are identical, the errors in the two channels follow the same probability distribution:
@@
In the first equality, the delta functions enforce the two conditions in Eq. (11), while the second equality carries over the summation using Eq. (13). The joint distribution cannot be factorized into a product of individual probability distributions, which means $s^{c}$ and $s^{t}$ are correlated random variables. This is exactly due to the shared error $E^{c}_{1}$ between $E^{c}$ and $E^{t}$ in Eq. (11).
@@
where $\{\sigma=\pm 1\},\{\tau=\pm 1\}$ are two sets of Ising spins that live on the lattice sites, and $l_{1},l_{2}$ are the two lattice sites at the two ends of the link $l$, similar to the labeling in Fig. 1(b). In this way, the trivial cycle $C$ can be separated into two parts: $C^{c}$ for the control block parameterized by $\{\sigma\}$, and $C^{t}$ for target code block parameterized by $\{\tau\}$, see Fig. 3. The stat-mech model can then be obtained via rewriting the summation of cycles using the Ising spins:
@@
where $S^{c,t}$ denote the stabilizers in the control and target code blocks, respectively. The disordered Hamiltonian is a 2D Ashkin-Teller (AT) model on the square lattice with random link couplings [42, 43, 44, 45, 46, 47]:
@@
where the coupling constants $K_{2}$ and $K_{4}$ are functions of $\tilde{p}$ (see Appendix A for their explicit forms), and $s^{c,t}_{l}$ are quenched random variables following the probability distribution in Eq. (III.3). We note that random AT model also arises from stat-mech mapping of toric code quantum memory under depolarizing noise [31] and both coherent and incoherent bit-flip noises [48, 34]. Our model in Eq. (17) differs from the existing models in the coupling constants and disorder distributions. Compared to RBIM for the memory case in Eq. (II.1), the two coupling constants in the random AT model satisfy
@@
for $\tilde{p}$ around the memory threshold (see Fig. 4). The relations between these coupling strengths have direct implications on the error thresholds of the control and target code blocks, as shown in the subsequent discussion.
Figure 3: Stat-mech mapping of two toric code blocks undergoing a tCNOT gate, where the probability distribution of $E^{c}$ and $E^{t}$ are correlated due to the tCNOT gate. Here the physical qubits are denoted by gray circles living on the lattice makred by solid lines. Bit-flip errors $E^{c/t}$ and $E^{c/t}\cdot C^{c/t}$ of the physical qubits in the control/target code blocks are marked by gray circles with colored fillings. Syndromes triggered by $E^{c/t}$ are marked by dots at the center of shaded plaquettes which host $Z$ stabilizers. Trivial cycles $C^{c/t}$ in control/target blocks are parameterized by two sets of Ising spins $\{\sigma\}$ and $\{\tau\}$, which are located at lattice sites of the dotted-line lattice, and are denoted by arrows in blue/green.
@@
The stat-mech model allows us to form a theoretical anticipation of the error thresholds of the two code blocks, even before carrying out the numerical simulation. In light of Eq. (4), the logical error rate for the control/target blocks now maps to the sum over $\overline{e^{-F_{DW}^{\sigma/\tau}}}$, where $F^{\sigma/\tau}_{DW}$ is the domain wall free energy of a global domain wall of the spins $\sigma/\tau$, and the disorder average is taken over both $E^{c}$ and $E^{t}$ (i.e. $s^{c}$ and $s^{t}$). Therefore, the error thresholds of the control and target code blocks, $p^{c}_{\text{c}}$ and $p^{t}_{\text{c}}$, correspond to the order-disorder phase transition of the $\sigma$ and $\tau$ spins, respectively. From Eq. (18), we expect that the target block will first reach the error threshold with increasing $\tilde{p}$. Since the critical point is within the ordered phase of $\{\sigma\}$ spins, we expect the coupling term between the $\{\sigma\}$ and $\{\tau\}$ spins to enhance the error threshold of the target block $p^{t}_{\text{c}}$. On the other hand, since $K_{2}<J$, the error threshold of the control layer $p^{c}_{\text{c}}$ should also be reduced compared to the case of quantum memory, $p_{\text{th}}\approx 0.109$.
III.4 Numerical investigations of the error thresholds
III.4.1 Monte Carlo simulation of the disordered AT model
@@
With the stat-mech model being constructed, we now conduct a numerical study to quantify the error thresholds of the two code blocks. Specifically, we perform a Monte Carlo (MC) simulation of the disordered AT model in Eq. (17), with disorders sampled from the distribution shown in Eq. (III.3). Unlike the conventional AT model, where $K_{2}$ and $K_{4}$ are independent parameters, our model depends on a single parameter $p$: $K_{2}$ and $K_{4}$ are both functions of $p$ (see Fig. 4). At each $p$, we first sample disorder configurations on all bonds $\{s_{l}^{t},s_{l}^{c}\}$ following Eq. (III.3). We then to equilibrate and sample the spin fields $\tau,\sigma$ with Hamiltonian described in Eq. (17) given quenched disorders $\{s_{l}^{t},s_{l}^{c}\}$. To ensure the MC simulation equilibrates, we introduce a hypothetical temperature $\beta$. The target distribution that we want to sample from is $e^{-\beta H}$ with $\beta=1$. We tune $\beta$ to effectively anneal the system from high temperature to target temperature $\beta=1$. We found this extra annealing equilibration stage to be crucial in reaching the equilibrium.
@@
To observe the order-to-disorder transition, we analyze the magnetization of the spin system at different $p$. As shown in Fig. 5, the magnetization $\tau$ and $\sigma$ spin fields $|M|$ both saturate to $1$ in the low noise limit ($p=0$). Since we start the MC from a random initialization, $|M|=1$ in the low $p$ limit affirms the convergence of the simulation. As the noise level increases, $\tau$ magnetization, corresponding to target code block, first starts decreasing and eventually vanishes. The magnetization of $\sigma$, corresponding to the control code block, tolerates a larger noise level before it starts to deviate from $|M|=1$. This observation is consistent with the fact that bit-flip errors propagate from the control code block to the target code block through the tCNOT operation, resulting in a lower threshold for the target code than for the control code.
@@
The major benefit of stat-mech mapping is the ability to leverage the established practice of finite-size scaling (FSS) to determine error thresholds in the infinite code distance limits. To obtain these thresholds, we need to collapse the magnetization of different system sizes $L=8,12,16$ into a unique scaling form [49]:
@@
As shown in Fig 6, we observe that the rescaled magnetization data achieve a remarkable collapse to universal functional forms around $\tilde{p}_{\text{c}}^{c}=0.052$ (or $p^{c}_{\text{c}}=0.099$, see Fig. 6(a)) and $\tilde{p}^{t}_{\text{c}}=0.042$ (or $p^{t}_{\text{c}}=0.080$, see Fig. 6(b)). These are compelling evidence of universal behavior in the vicinity of these critical points. Therefore, these critical points can be interpreted as upper bounds on decoding thresholds for control ($\sigma$) and target ($\tau$) code blocks in the large code distance limit. Compared to threshold of quantum memory, $p_{\text{th}}=0.109$,the thresholds of the control and the target blocks are lowered by $\frac{0.109-0.099}{0.109}=9.2\%$ and $\frac{0.109-0.080}{0.109}=26.6\%$, respectively. These results confirm our theoretical expectation at the end of Sec. III.3 regarding the hierarchy of the reduced thresholds. At the same time, our results establish that tCNOT gates do not lower the thresholds to a devastatingly degree.
Our finite size scaling results also highlight the benefit of correlated decoding. If one decodes the target block separately, the detectors of the target block effectively experience three consecutive bit-flip channels of $\mathcal{N}_{\tilde{p}}$. The error threshold of such a channel is the solution to
@@
Figure 7: Logical error rate of a surface under same noise model using then most likely error decoder. The crossing point of curves for different system sizes indicates a decoding threshold. The gray lines represent the phase transition point predicted by Monte Carlo simulation of the random AT model.
III.4.2 Direct simulation of error correction protocol
We cross-check the thresholds predicted by the MC simulation of the AT model with those observed in a practical error correction protocol. Since the error thresholds for a planar surface code and the toric code are expected to be the same [8], we opt to simulate a rotated surface code for convenience. We use stim package [41] to simulate and sample syndromes from the same logical circuit shown in Fig. 2(a) under the noise model described in Fig. 2(b). We then decode the sampled syndromes using the most likely error (MLE) decoder [24] and record the resulting logical error rates (LERs). Although the MLE decoder is not scalable and thus not favorable in practical QEC, it serves as a valuable tool for validating our theoretical predictions and identifying performance bounds. As shown in Fig. 7, the LERs of the control and target code with different code distances each cross at one particular error rate, which is the observed decoding threshold. The crossings of LER for both control and target codes match the critical points obtained from FSS analysis of magnetization data from MC simulation of stat-mech mapped disordered AT model. Below the decoding threshold, increasing code distance suppresses logical errors.
IV Toric codes under tCNOT gate with syndrome errors

### v3

@@ 3.1 Persistent bit-flip error model
Therefore, we consider an error model that captures the spread of errors, in which bit-flip errors occur both before and after the tCNOT gate. To directly compare with the case of quantum memory, we evenly split the error rate of channel $\mathcal{N}_{p}$ into two bit-flip channels $\mathcal{N}_{\tilde{p}}$ before and after the tCNOT gate:
@@
We will subsequently refer to such an error model as the persistent bit-flip error [^3] When referring to the bit-flip error without the word “persistent”, we assume the bit-flip channel in Eq. (6) occurs only after the tCNOT gate, which is the case in the noise model we study in Sec. 4. .
@@
3.2 Detectors of toric codes across a tCNOT gate
The main challenge in applying stat-mech mapping to logical circuits lies in the need to track dynamical error propagation through gates. To this end, we import the quantum circuit simulation notion of a detector [41] and integrate it into the stat-mech mapping procedure. In our case, a detector associated with a stabilizer measurement outcome is a product of that outcome with earlier stabilizer measurement outcomes that back-tracks the flow of the measured Pauli operator in the circuit. Without errors, each detector’s value is deterministically 1. When an error $E$ occurs in the circuit, a detector can be triggered, causing its value to flip to -1. The collection of detector values, $\{d\}$, serves as the syndrome. Therefore, once the detectors and its relation with every error $E$ in the error model is specified, so will the cycles $C$.
In the context of a logical circuit with a tCNOT gate, the spacetime detectors are fixed in the following way [24, 26, 27, 28, 29]. Since we consider only bit-flip errors, we backtrack the spread of the measured $Z$ stabilizers. In this way, the spacetime detectors between the two adjacent syndrome extraction rounds are the ones shown in Fig. 2(a): the spacetime detector $d^{c}$ of the control code block is the product of the measurement outcomes $M^{c}_{1/2}$ of the $Z$ stabilizers before/after the tCNOT gate:
@@ 3.2 Detectors of toric codes across a tCNOT gate
where $M^{t}_{1,2}$ are outcomes of the $Z$ stabilizer measurements before and after the tCNOT gate. The superscripts denote the control code ($c$) and the target code ($t$), respectively (alternatively, one can also choose $d^{t}=M^{c}_{2}M^{t}_{1}M^{t}_{2}$). A visualization of the detectors is shown in Fig. 2(a). The spacetime detectors directly incorporate both error-propagation mechanisms arising from the tCNOT gate. Namely, a bit-flip error that occurs in the error channel $\mathcal{N}_{\tilde{p}}$ before the tCNOT gate in the control block now triggers both $d^{c}$ and $d^{t}$.
Figure 2: (a) The implementation of the tCNOT gate between two rounds of syndrome extractions. The black dots at the centers of the shaded plaquettes represent measurements of the $Z$ stabilizers defined on these plaquettes. The choices of spacetime detectors $d^{c}$ and $d^{t}$ for the control and target code blocks between the two rounds of SE are marked in blue and green, respectively. Due to the tCNOT gate, the spacetime detector of the target block, $d^{t}$, extends to the control block, whose value is obtained from multiplying three $Z$ stabilizer measurement outcomes $d^{t}=M^{c}_{1}M^{t}_{1}M^{t}_{2}$. To avoid visual overlap, we sketch the spacetime detectors $d^{c}$ and $d^{t}$ in two adjacent plaquettes that host $Z$ stabilizers. (b) The logical circuit with tCNOT gate, persistent bit-flip noise channels $\mathcal{N}_{\tilde{p}}$ and two rounds of perfect syndrome extractions (SE), during which the weight-4 $Z$ stabilizers are measured.
3.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
With the detectors and error model fixed, we now construct the stat-mech model of the two toric code blocks across the tCNOT gate. We denote the bit-flip errors that occur during the four error channels $\mathcal{N}_{\tilde{p}}$ in Fig. 2(a) as $E^{c,t}_{1,2}$, where $c/t$ indicates the control/target code blocks and $1/2$ denotes the error before/after the tCNOT gate. Given the form of the spacetime detectors $d^{c/t}$, the errors that trigger the spacetime detectors in the control and target blocks are
@@ 3.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
We focus on a pair of physical qubits in the control and target blocks. Since the error channels $\mathcal{N}_{\tilde{p}}$ are identical, the errors in the two channels follow the same probability distribution:
@@
In the first equality, the delta functions enforce the two conditions in Eq. (11), while the second equality carries over the summation using Eq. (13). The joint distribution cannot be factored into a product of individual probability distributions, so $s^{c}$ and $s^{t}$ are correlated random variables. This is exactly due to the shared error $E^{c}_{1}$ between $E^{c}$ and $E^{t}$ in Eq. (11).
@@
where $\{\sigma=\pm 1\},\{\tau=\pm 1\}$ are two sets of Ising spins that live on the lattice sites, and $l_{1},l_{2}$ are the two lattice sites at the two ends of the link $l$, similar to the labeling in Fig. 1(b). In this way, the trivial cycle $C$ can be separated into two parts: $C^{c}$ for the control block parameterized by $\{\sigma\}$, and $C^{t}$ for the target code block parameterized by $\{\tau\}$, see Fig. 3. The stat-mech model can then be obtained via rewriting the summation of cycles using the Ising spins:
@@
where $S^{c,t}$ denotes the stabilizers in the control and target code blocks, respectively. The disordered Hamiltonian is a 2D Ashkin-Teller (AT) model on the square lattice with random link couplings [42, 43, 44, 45, 46, 47]:
@@
where the coupling constants $K_{2}$ and $K_{4}$ are functions of $\tilde{p}$ (see Appendix A for their explicit forms), and $s^{c,t}_{l}$ are quenched random variables following the probability distribution in Eq. (3.3). We note that random AT model also arises from stat-mech mapping of toric code quantum memory under depolarizing noise [31] and both coherent and incoherent bit-flip noises [48, 34]. Our model in Eq. (17) differs from the existing models in the coupling constants and disorder distributions. Compared to RBIM for the memory case in Eq. (2.1), the two coupling constants in the random AT model satisfy
@@
for $\tilde{p}$ around the memory threshold (see Fig. 4). The relations between these coupling strengths have direct implications for the error thresholds of the control and target code blocks, as shown in the subsequent discussion.
Figure 3: Stat-mech mapping of two toric code blocks undergoing a tCNOT gate, where the probability distribution of $E^{c}$ and $E^{t}$ is correlated due to the tCNOT gate. Here, the physical qubits are denoted by gray circles living on the lattice marked by solid lines. Bit-flip errors $E^{c/t}$ and $E^{c/t}\cdot C^{c/t}$ of the physical qubits in the control/target code blocks are marked by gray circles with colored fillings. Syndromes triggered by $E^{c/t}$ are marked by dots at the center of shaded plaquettes, which host $Z$ stabilizers. Trivial cycles $C^{c/t}$ in control/target blocks are parameterized by two sets of Ising spins $\{\sigma\}$ and $\{\tau\}$, which are located at lattice sites of the dotted-line lattice, and are denoted by arrows in blue/green.
@@
The stat-mech model allows us to predict the error thresholds of the two code blocks, even before performing the numerical simulation. In light of Eq. (4), the logical error rate for the control/target blocks now maps to the sum over $\overline{e^{-F_{DW}^{\sigma/\tau}}}$, where $F^{\sigma/\tau}_{DW}$ is the domain wall free energy of a global domain wall of the spins $\sigma/\tau$, and the disorder average is taken over both $E^{c}$ and $E^{t}$ (i.e. $s^{c}$ and $s^{t}$). Therefore, the error thresholds of the control and target code blocks, $p^{c}_{\text{c}}$ and $p^{t}_{\text{c}}$, correspond to the order-disorder phase transition of the $\sigma$ and $\tau$ spins, respectively. From Eq. (18), we expect that the target block will first reach the error threshold with increasing $\tilde{p}$. Since the critical point is within the ordered phase of $\{\sigma\}$ spins, we expect the coupling term between the $\{\sigma\}$ and $\{\tau\}$ spins to enhance the error threshold of the target block $p^{t}_{\text{c}}$. On the other hand, since $K_{2}<J$, the error threshold of the control layer $p^{c}_{\text{c}}$ should also be reduced compared to the case of quantum memory, $p_{\text{th}}\approx 0.109$.
3.4 Numerical investigations of the error thresholds
3.4.1 Monte Carlo simulation of the disordered AT model
@@
With the stat-mech model being constructed, we now conduct a numerical study to quantify the error thresholds of the two code blocks. Specifically, we perform a Monte Carlo (MC) simulation of the disordered AT model in Eq. (17), with disorders sampled from the distribution shown in Eq. (3.3). Unlike the conventional AT model, where $K_{2}$ and $K_{4}$ are independent parameters, our model depends on a single parameter $p$: $K_{2}$ and $K_{4}$ are both functions of $p$ (see Fig. 4). At each $p$, we first sample disorder configurations on all bonds $\{s_{l}^{t},s_{l}^{c}\}$ following Eq. (3.3). We then to equilibrate and sample the spin fields $\tau,\sigma$ with Hamiltonian described in Eq. (17) given quenched disorders $\{s_{l}^{t},s_{l}^{c}\}$. To ensure the MC simulation equilibrates, we introduce a hypothetical temperature $\beta$. The target distribution that we want to sample from is $e^{-\beta H}$ with $\beta=1$. We tune $\beta$ to anneal the system from a high temperature to the target temperature $\beta=1$. We found this additional annealing equilibration stage to be crucial for reaching equilibrium.
@@
To observe the order-to-disorder transition, we analyze the magnetization of the spin system as a function of $p$. As shown in Fig. 5, the magnetization $\tau$ and $\sigma$ spin fields $|M|$ both saturate to $1$ in the low noise limit ($p=0$). Since we start the MC from a random initialization, $|M|=1$ in the low-$p$ limit confirms the convergence of the simulation. As the noise level increases, $\tau$ magnetization, corresponding to the target code block, first starts decreasing and eventually vanishes. The magnetization of $\sigma$, corresponding to the control code block, tolerates a larger noise level before it starts to deviate from $|M|=1$. This observation is consistent with the fact that bit-flip errors propagate from the control code block to the target code block through the tCNOT operation, resulting in a lower threshold for the target code than for the control code.
@@
The major benefit of stat-mech mapping is the ability to leverage the established practice of finite-size scaling (FSS) to determine error thresholds in the infinite-code-distance limit. To obtain these thresholds, we need to collapse the magnetization of different system sizes $L=8,12,16$ into a unique scaling form [49]:
@@
As shown in Fig 6, we observe that the rescaled magnetization data achieve a remarkable collapse to universal functional forms around $\tilde{p}_{\text{c}}^{c}=0.052$ (or $p^{c}_{\text{c}}=0.099$, see Fig. 6(a)) and $\tilde{p}^{t}_{\text{c}}=0.042$ (or $p^{t}_{\text{c}}=0.080$, see Fig. 6(b)). These are compelling evidence of universal behavior in the vicinity of these critical points. Therefore, these critical points can be interpreted as upper bounds on the decoding thresholds for control ($\sigma$) and target ($\tau$) code blocks in the large-code-distance limit. Compared to the threshold of quantum memory, $p_{\text{th}}=0.109$, the thresholds of the control and the target blocks are lowered by $\frac{0.109-0.099}{0.109}=9.2\%$ and $\frac{0.109-0.080}{0.109}=26.6\%$, respectively. These results confirm our theoretical expectation at the end of Sec. 3.3 regarding the hierarchy of the reduced thresholds. At the same time, our results establish that tCNOT gates do not lower the thresholds to a devastating degree.
Our finite-size scaling results also highlight the benefit of correlated decoding. If one decodes the target block separately, the detectors of the target block effectively experience three consecutive bit-flip channels of $\mathcal{N}_{\tilde{p}}$. The error threshold of such a channel is the solution to
@@
Figure 7: Logical error rate of a surface code under the same noise model using the most likely error decoder. The crossing point of curves for different system sizes indicates a decoding threshold. The gray lines represent the phase transition points predicted by Monte Carlo simulation of the random AT model.
3.4.2 Direct simulation of error correction protocol
We cross-check the thresholds predicted by the MC simulation of the AT model against those observed in a practical error-correction protocol. Since the error thresholds for a planar surface code and the toric code are expected to be the same [8], we opt to simulate a rotated surface code for convenience. We use stim package [41] to simulate and sample syndromes from the same logical circuit shown in Fig. 2(a) under the noise model described in Fig. 2(b). We then decode the sampled syndromes using the most likely error (MLE) decoder [24] and record the resulting logical error rates (LERs). Although the MLE decoder is not scalable and thus not suitable for practical QEC, it remains a valuable tool for validating our theoretical predictions and identifying performance bounds. As shown in Fig. 7, the LERs of the control and target code with different code distances each cross at one particular error rate, which is the observed decoding threshold. The crossings of LER for both control and target codes match the critical points obtained from FSS analysis of magnetization data from MC simulations of a stat-mech-mapped disordered AT model. Below the decoding threshold, increasing code distance suppresses logical errors.
4 Toric codes under tCNOT gate with syndrome errors